OpenStudy (anonymous):

If a copper wire has a resistance of 19.2 Ω at a temperature of 22.0 oC, what resistance does it have at 79.7 oC? (Neglect any change in length or cross-sectional area resulting from the change in temperature.)

OpenStudy (anonymous):

ΔR=αR0ΔT. the law is for resistivity, but because length and area will be the same then you can convert resistivity to resistance. alpha is a constant

OpenStudy (anonymous):

http://hyperphysics.phy-astr.gsu.edu/hbase/electric/restmp.html if you used the constant in the above link \[r-19.2 = \alpha \times 19.2 \times ( 57.7)\] the alpha should be given

OpenStudy (anonymous):

I guess I am still confused, I don't think alpha is given

OpenStudy (kropot72):

http://www.cirris.com/learning-center/testing-guidelines/special-topics/177-temperature-coefficient-of-copper The information here gives you the temperature coefficient of resistance for copper, and gives examples of calculations that show you how to solve your question.

OpenStudy (anonymous):

Any way alpha should mention in previous problem or in a table in your book.